OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. 2 – 8x + 3x when x = 5 A. –53 B. –23 C. –15 D. 57

OpenStudy (anonymous):

no b

OpenStudy (anonymous):

my bad its B -40+15= -26+2= -23